OpenStudy (chris215):

PLEASE HELP: Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y=7/(x^2+10)

OpenStudy (anonymous):

you must look at the formula and see (at least) to steps. Well, if I look at it, I see a complicated first step: 7/x² and a simple second one: adding 10 to a number.

OpenStudy (anonymous):

Go to google, and look up 'Tiger Algebra'. Enter your problem and it can solve it.

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

\[f(x)=x+10\]

OpenStudy (lochana):

g(x) = x^2 ; f(x) = 7/(x +10)

OpenStudy (anonymous):

Now if you have to calculate f(g(x)), this means: put the result of g, which is 7/x² in f. This gives the formula you had: 7/x² + 10.

OpenStudy (anonymous):

See? g(x)=7/x² has taken the place of x in function f.

OpenStudy (anonymous):

WTF

OpenStudy (anonymous):

???????????????????????????????????????????? !!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

nvm

OpenStudy (lochana):

@chris215 see my answer, I've answered to this question before

OpenStudy (chris215):

thank you for the help!!